40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  PSL  Not logged in ELOG logo
Entry  Tue Apr 10 00:50:56 2012, frank, tara, DailyProgress, optic, bad PBS 
    Reply  Wed Apr 11 02:44:29 2012, frank, tara, DailyProgress, optic, bad PBS beat_2012_04_10.png
       Reply  Wed Apr 11 20:30:50 2012, frank, tara, DailyProgress, NoiseBudget, Electronic noise verification beat_2012_04_11.png
          Reply  Thu Apr 12 22:22:12 2012, frank, tara, DailyProgress, NoiseBudget, Electronic noise + shot noise verification 
Message ID: 906     Entry time: Wed Apr 11 02:44:29 2012     In reply to: 905     Reply to this: 909
Author: frank, tara 
Type: DailyProgress 
Category: optic 
Subject: bad PBS 

Update, beat measurement after several optics replacement. Peaks around 10 Hz, 35 Hz show up this time.

    Optics that we replaced are:

  •  Beam splitter that divides the beam to ACAV and RCAV path. Now the new BS is 1" cube for large beam
  • Fixed the orientation of mirrors on the periscope and the turning mirror for ACAV RFPD. A few of them were flipped back causing stray light in the beam path.
  • Two of the mirrors on periscope had transmission of 2% or so, we replaced with a high reflective ones (0.1% transmission).

The problem with the curve mirror from last entry has not been fixed yet. It turns out that the mirror we borrow from 40m is worse than the one we have (surface is more milky), so we leave the original mirror as it is.

Note: The beat measurement was done when the air springs were inactive. Noise at high frequency goes down a bit.

The power input to each cavity is 1mW, setup on PLL is 1kHz input range, with gain = 200.

beat_2012_04_10.png

ELOG V3.1.3-